# Radio Morphing: Fast computation of inclined air shower radio emission

### v1

@@ 4.4 Computation time
Figure 9: (Left) Peak-time delay at the antenna level between Radio Morphing and ZHAireS simulations. (Right) Mean CPU time (square) and standard deviation (error bars) of Radio Morphing simulations as a function of the shower zenith angle. For each shower, the radio emission was simulated at $176$ antenna positions with a thinning of $10^{-5}$.
Eventually, we evaluated the CPU time of Radio Morphing simulations, since the primary objective of the tool is to simulate air shower radio signals for future large-scale radio experiments such as GRAND or the Pierre Auger Observatory. In the right-hand panel of Fig. 9, we show the mean CPU time of Radio Morphing as a function of the shower zenith angle from the set of $4500$ showers generated with Radio Morphing to test its performances. Each simulation corresponds to a thinning of $10^{-5}$ and was run for $176$ antennas with a primary energy between $10^{17}\,\rm eV$ and $4\cdot 10^{18}\,\rm eV$. It can be seen that the computation time increases with the zenith angle, which is expected, as more reference planes are used for inclined showers. Still, Radio Morphing provides a computation time of only a few tens of seconds, independently of the shower zenith angle and primary energy. For a quantitative comparison, simulations with the same thinning, primary energy, and number of antennas, run with ZHAireS or CoREAS, are generally expected to take $\sim 1-2$ days of CPU time (with AMD EPYC 7302 16-Core Processor) for $\theta=60^{\circ}$ and around 1 week for $\theta=85^{\circ}$. This leads to an improvement of more than 4 orders of magnitude, making Radio Morphing a suitable tool to run a massive number of simulations in a short amount of time, in preparation for future large-scale radio detectors.
@@ 5 Conclusion
We presented the latest version of Radio Morphing, a semi-analytical method that allows for a fast computation of any cosmic-ray induced air shower with zenith angle $\theta>60^{\circ}$, at any desired position in space, from the simulation data of a few reference ZHAireS showers at given positions. This new version now incorporates shower-to-shower fluctuations, refined scaling laws, interpolation method, and modeling of charge-excess emission. Radio Morphing performances were evaluated by comparison to ZHAireS simulations ran for the same parameters. For most zenith angles, we find mean relative differences in the full band peak amplitude around $5\%$, indicating only a minor bias of the method, mainly for low-energy showers with an azimuthal direction close to the magnetic North. The corresponding standard deviations on raw traces were evaluated to be below $17\%$ independently of the zenith angle, i.e., below the calibration error reached by current radio detectors. The standard deviations were even reduced to $15\%$ when applying a $3\sigma$ trigger threshold and below $\sim 10\%$ when working in the $30-80]\,\rm MHz$ band. These performances were coupled with a timing accuracy of less than $5\,\rm ns$ and a time computation reduced by more than four orders of magnitude compared to standard Monte Carlo simulations with primary energy above $10^{17}\,\rm eV$. This makes Radio Morphing a powerful and accurate tool for simulating a huge number of air shower simulations in the view of future large-scale radio detectors such as GRAND, SKA, or the radio upgrade of the Pierre Auger Observatory. Further improvements would include the correction of the bias with the shower azimuth angle, an implementation of radio emission scaling with the Earth’s magnetic field and the incorporation of other primaries such as neutrinos. Radio Morphing open source code is written in python and now publicly available at https://github.com/simonchiche/RadioMorphing.

### v2

@@ 4.4 Computation time
Figure 9: Peak-time difference at the antenna level between Radio Morphing and ZHAireS simulations.
Figure 10: Mean CPU time (square) and standard deviation (error bars) of Radio Morphing simulations as a function of the shower zenith angle. For each shower, the radio emission was simulated at $176$ antenna positions with a thinning of $10^{-5}$.
Eventually, we evaluated the CPU time of Radio Morphing simulations, since the primary objective of the tool is to simulate air shower radio signals for future large-scale radio experiments such as GRAND or the Pierre Auger Observatory. In Fig. 10, we show the mean CPU time of Radio Morphing as a function of the shower zenith angle from the set of $4500$ showers generated with Radio Morphing to test its performances. Each simulation corresponds to a thinning of $10^{-5}$ and was run for $176$ antennas with a primary energy between $10^{17}\,\rm eV$ and $4\cdot 10^{18}\,\rm eV$. It can be seen that the computation time increases with the zenith angle, which is expected, as more reference planes are used for inclined showers. Still, Radio Morphing provides a computation time of only a few tens of seconds, independently of the shower zenith angle and primary energy. For a quantitative comparison, simulations with the same thinning, primary energy, and number of antennas, run with ZHAireS or CoREAS, are generally expected to take $\sim 1-2$ days of CPU time (with AMD EPYC 7302 16-Core Processor) for $\theta=60^{\circ}$ and around 1 week for $\theta=85^{\circ}$. This leads to an improvement of more than 4 orders of magnitude, making Radio Morphing a suitable tool to run a massive number of simulations in a short amount of time, in preparation for future large-scale radio detectors.
@@ 5 Conclusion
We presented the latest version of Radio Morphing, a semi-analytical method that allows for a fast computation of any cosmic-ray induced air shower with zenith angle $\theta>60^{\circ}$, at any desired position in space, from the simulation data of a few reference ZHAireS showers at given positions. This new version now incorporates shower-to-shower fluctuations, refined scaling laws, interpolation method, and modeling of charge-excess emission. Radio Morphing performances were evaluated by comparison to ZHAireS simulations ran for the same parameters. For most zenith angles, we find mean relative differences in the full band peak amplitude around $5\%$, indicating only a minor bias of the method, mainly for low-energy showers with an azimuthal direction close to the magnetic North. The corresponding standard deviations on raw traces were evaluated to be below $17\%$ independently of the zenith angle, i.e., below the calibration error reached by current radio detectors. The standard deviations were even reduced to $15\%$ when applying a $3\sigma$ trigger threshold and below $\sim 10\%$ when working in the $30-80]\,\rm MHz$ band. These performances were coupled with a timing accuracy of less than $5\,\rm ns$ and a time computation reduced by more than four orders of magnitude compared to standard Monte Carlo simulations with primary energy above $10^{17}\,\rm eV$. This makes Radio Morphing a powerful and accurate tool for simulating a huge number of air shower simulations in the view of future large-scale radio detectors such as GRAND, SKA, or the radio upgrade of the Pierre Auger Observatory. We note that the current version of Radio Morphing is tuned to the magnetic field configuration of Dunhuang (China). Nevertheless, the method is intrinsically generic and remains applicable to other magnetic field configurations. Extending it to a different site would only require adapting the reference simulations, which could be performed upon request. Further improvements would include the correction of the bias with the shower azimuth angle, an implementation of radio emission scaling with the Earth’s magnetic field and the incorporation of other primaries such as neutrinos. Radio Morphing open source code is written in python and now publicly available at https://github.com/simonchiche/RadioMorphing.